i really need help i am an at home student so its hard to do some of the math
(4-6i)-2(5+2i)-i

Answer by sudhanshu_kmr(1152)


(4-6i)-2(5+2i)-i
= (4-6i)-10 -4i -i
= 4-6i -10 -4i -i
= -6 - 11i